OpenStudy (anonymous):

olie uses 2 tomatoes every day to prepare her salad. If n equals the number of tomatoes Jolie had before she made her salad and c equals the number of tomatoes left after the salad is made, which equation represents the number of tomatoes Jolie has after she made her salad?

OpenStudy (anonymous):

n-2=c? But I don't know how many days she made a salad. is it asking for just that day?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i dont think thats right

OpenStudy (anonymous):

it says 2 tomatoes for her salad. ok well you had n amount before you made it. Well say you had 7. You used 2 so now you have 5. replace the numbers for letters. n-2=c

OpenStudy (anonymous):

would you like me to invite a friend to help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok here is someone who helped me out alot. @student_basil can you help desiree412?

OpenStudy (texaschic101):

Is this multiple choice ?

OpenStudy (anonymous):

yes n=2-c C=n-2 c=2-n n=c-2

OpenStudy (anonymous):

c=n-2

OpenStudy (texaschic101):

C = n - 2

OpenStudy (anonymous):

that is basically what i've been saying but just did not put the c first...

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

thanks for the medals

OpenStudy (texaschic101):

give credit where credit is due :)

OpenStudy (anonymous):

lol